If you triple the voltage across a resistor while at the same time cutting its resistance to one-fifth its original value, what happens to the current in the resistor?


If you triple the voltage across a resistor while at the same time cutting its resistance to one-fifth its original value, what happens to the current in the resistor?

it increases by 15 times
it doubles
it increases by five times
you can't tell from the data given
Answer #1

I = V/R = 3V/R/5 = 15 V/R


so it increases by 15 times
Answer #2

as per ohms law V = iR

if V is tripled and R is reduced to 1/5th

i = v/R

i' = 3V/R/5 = 15 i

it increases by 15 times
Answer #3

Increaswes by 15 times.

I = V / R


I' = (5V) / (R/5) = 15 V / R = 15 I